CH/Agen/92/05

Agenda

Charissa Group Meeting

To be held on: 7 October 1992, 11.30 am To be held at: Surrey, Jackson Room (16AC22)

Distribution: WNC,NMC,RAC,BRF,GJG,RAH,JSL,ACM,VP,WDMR,GT,DLW

1.  Minutes of Meeting of 24 August 1992
2.  Matters Arising/ Actions
3.  Update on NSC/SERC matters  (DLW)
4.  Report on trip to the ANU  (WDMR)
5.  Progress on scintillators  (NMC)
6.  Progress on silicon detectors  (BRF)
7.  Progress on gas supply  (RAC)
8.  Progress on chamber & detector assemblies  (JSL)
9.  Progress on electronics & data acquisition  (RAH)
10.  Developments on hybrid detectors  (RAC)
11.  Developments on event manager & GEC emulator  (RAH,VP)
12.  Update on Recent Analysis  (BRF,WDMR,

WNC,DLW)

13.  Discussion of Future Plans

(NSF runs, Berkeley, ANU plus others)

14.  Any other business
15.  Date of next meeting

W.N. Catford (Secretary) 23 September 1992